RC meeting and Run Plan, Thursday Sept. 21, 2023
Return to the RC Meeting Calendar
- 1 RC Meeting Information
- 2 Useful Links
- 3 Procedures
- 4 Status Update
RC Meeting Information
Time:at 3:30pm today
- Location: Counting house conference room
- ZoomGov Meeting Connection Information
Hall A call-in instructions:
- Call RC first.
- Techs on site 7-4:30 Mon-Th, 7-11AM Fr. RC will call Hall A work coordinator (Zach) during normal working hours. After hours, RC calls Tech on-call.
- After 9pm everything non-emergency needs to be deferred. What is emergency and what's not? Call RC. RC talks with Mark to decide.
Hall Access procedure:
- Access check list (binder on the counting room desk) needs to be filled out before making an access.
- Scan (take pictures with phone) and post it to the logbook.
- Any work inside the SBS, BB boundary need to have PS locked out -- requires two QEWs to throw the big switch. Exception: field measurements.
New instruction for shift crew when changing target:
- Beam restoration instructions have been updated. Printed version is in the counting room.
- When we change target, make sure the SOFs of Target BPMs are set to the correct values.
- SOFs for different targets: Needs an update after restoring beam tonight (Note: SOFs for "No Target" has been updated and they are same as the He3 target values)
During the maintenance day yesterday, we accomplished the SBS magnet acid flush, carbon hole target change and fixing issues with different detectors. After all the work, hall was taken to controlled access around 6pm. During the controlled access, HCAL hardware is checked again to make sure that it's functioning well. We also changed half wave plate from OUT to IN. Around 7pm, hall was taken to beam permit and we got beam to the hall after 9 pm. Beam recovery procedure wall followed thoroughly, and it was completed around early owl shift. New carbon hole center is found at (0.8,-1.1) https://logbooks.jlab.org/entry/4184156 along with the previous SOFs. After that MCC struggled to send the CW to the reference cell during the whole owl shift. As planned, Moeller measurement was started from the early day shift today, but there was issue with Moller dipole which requires ~4hr to fix by DC power.
- MCC struggled to steer beam to Reference cell during OWL shift (target ion chambers screaming)
- Work required in NL transfer https://logbooks.jlab.org/entry/4184580 and https://tasklists.jlab.org/atlis/tasks/109934
- Maintenance day (Tuesday 09/26/2023)
Charge collected (polarization corrected):
- 0 C
Hall A Subsystems Status
- Polarization: 26.3% (12:08 pm)
Known Issues for shiftcrews
- FSD engaged due to cooling jet limit ==> found out that it's intermittent only but needs attention
- SBS GEMs made DAQ unstable ==> need to change NoSBSGEMs config
- Hodoscope alarm going off even though HV is okay ==> Needs attention
- HCAL/BBCALThreshold GUI readback ==> Ellen has RPI and needs to plug back during the hall access
- Looks very good, fully staffed for the next week.
- Complete the Moeller measurement after fixing the Moeller dipole by DC power
- Carbon-hole target center check: with new SOFs in their script
- Reference cell H2 data taking or Production data with He3 (??)
List of items in case target isn’t ready:
- Take data on reference cell at 10 atm H2 (portions that add up to 1 day)
- BigBite Sieve slit => only during day (Techs needed)
- with 8 carbon foils and take optics for 1 hr at 40 uA (Sean/Andrew - confirm time)
- 1 hour for single foil (magnets off) => Chance BBcal HV setting to appropriate values
- BigBite/single foil location reconstruction (~20 mins?) - BigBite + HCAL detector experts on call consulting analysis experts like Andrew/Sean (magnet on)